# LangForce: Bayesian Decomposition of Vision Language Action Models via Latent Action Queries

### v4

@@ 7 Conclusion
In this work, we identified a critical pathology in standard VLA training: the vision shortcut, where models learn to ignore language instructions in goal-driven datasets, leading to a collapse of the mutual information between instructions and actions. This phenomenon results in catastrophic failure when the model faces ambiguous scenarios or out-of-distribution environments where visual shortcuts no longer hold. To address this, we introduced LangForce, a novel framework that formalizes VLA learning through a Bayesian lens. By decomposing the policy into a vision-only prior and a language-conditioned posterior, we derived a training objective that explicitly maximizes the Log-Likelihood Ratio (LLR), effectively optimizing the Pointwise Mutual Information (PMI) between language and action. We instantiated this decomposition using a unified architecture with Latent Action Queries and a dual-branch training strategy, which incurs no additional computational overhead during inference. Our extensive experiments on SimplerEnv and RoboCasa demonstrate that LangForce effectively mitigates the vision shortcut problem. It significantly outperforms baselines under various conditions; in out-of-distribution simulation environments, where standard methods fail completely, LangForce achieves robust generalization. These results highlight the importance of introducing structural inductive biases in learning robust, generalizable robot policies. We hope our work inspires further research into the information-theoretic foundations of embodied AI, pushing towards robots that truly understand and reliably execute human intent.
@@
For future work, we intend to provide a more comprehensive empirical evaluation in subsequent versions of this manuscript. We plan to extend our experimental validation to include the RoboTwin2 (Chen et al., 2025) and BEHAVIOR-1K (Li et al., 2024a) benchmarks, alongside real-world robot experiments. We also aim to scale our framework to larger foundation models, such as conducting experiments with LangForce on Qwen3VL-8B. Additionally, we plan to include more extensive ablation studies to further dissect the contribution of each component and hyperparameter within our Bayesian decomposition framework, thereby rendering the analysis more complete.
@@ Appendix A Probabilistic Analysis of the Vision Shortcut
Consequently, as $H(\ell\mid v)\to 0$, the upper bound on the mutual information vanishes, forcing $I(\ell;a\mid v)\to 0$ (Xu et al., 2025). In this regime, $\mathbb{E}[\log\mathcal{R}]\approx 0$, rendering the specific LLR objective uninformative on average for samples drawn from such biased distributions. This theoretical hypothesis is strongly supported by our empirical results in Appendix LABEL:app:robocasa, where a Vision-Only policy achieves a high success rate of 44.7% on RoboCasa, nearly matching the instruction-following baseline (47.8%). This confirms that for a significant portion of the data, $v$ indeed suffices to infer the task.

### v5

@@ 7 Conclusion
In this work, we identified a critical pathology in standard VLA training: the vision shortcut, where models learn to ignore language instructions in goal-driven datasets, leading to a collapse of the mutual information between instructions and actions. This phenomenon results in catastrophic failure when the model faces ambiguous scenarios or out-of-distribution environments where visual shortcuts no longer hold. To address this, we introduced LangForce, a novel framework that formalizes VLA learning through a Bayesian lens. By decomposing the policy into a vision-only prior and a language-conditioned posterior, we derived a training objective that explicitly maximizes the Log-Likelihood Ratio (LLR), effectively optimizing the Pointwise Mutual Information (PMI) between language and action. We instantiated this decomposition using a unified architecture with Latent Action Queries and a dual-branch training strategy, which incurs no additional computational overhead during inference. Our extensive experiments on SimplerEnv, LIBERO and RoboCasa demonstrate that LangForce effectively mitigates the vision shortcut problem. It significantly outperforms baselines under various conditions; in out-of-distribution simulation environments, where standard methods fail completely, LangForce achieves robust generalization. These results highlight the importance of introducing structural inductive biases in learning robust, generalizable robot policies. We hope our work inspires further research into the information-theoretic foundations of embodied AI, pushing towards robots that truly understand and reliably execute human intent.
@@
For future work, we intend to provide a more comprehensive empirical evaluation in subsequent versions of this manuscript. We plan to extend our experimental validation to include the RoboTwin2 (Chen et al., 2025) and BEHAVIOR-1K (Li et al., 2024a) benchmarks, alongside real-world robot experiments. We also aim to scale our framework to larger foundation models, such as conducting experiments with LangForce on Qwen3VL-8B or Qwen3VL-30B-A3B (Bai et al., 2025).
@@ Appendix A Probabilistic Analysis of the Vision Shortcut
Consequently, as $H(\ell\mid v)\to 0$, the upper bound on the mutual information vanishes, forcing $I(\ell;a\mid v)\to 0$ (Xu et al., 2025). In this regime, $\mathbb{E}[\log\mathcal{R}]\approx 0$, rendering the specific LLR objective uninformative on average for samples drawn from such biased distributions. This theoretical hypothesis is strongly supported by our empirical results in Section 2.1, where a Vision-Only policy achieves a high success rate of 44.7% on RoboCasa, nearly matching the instruction-following baseline (47.8%). This confirms that for a significant portion of the data, $v$ indeed suffices to infer the task.
@@ Appendix B Derivation of the LLR Objective
Appendix C Additional Experiments
C.1 Ablation on $\lambda$ and $\beta$ in Loss Function
Table 6: Ablation study on $\beta$ in Eq. 9. All experiments are based on the Qwen3-VL-4B backbone.
$\lambda$ Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average 0 86.4 64.6 26.0 76.0 63.3 0.1 84.2 65.6 27.3 81.9 64.8 0.3 89.6 63.8 33.3 79.2 66.5 0.5 81.2 61.5 25.8 87.7 64.1
Table 7: Ablation study on $\beta$ in Eq. 9. All experiments are based on the Qwen3-VL-4B backbone.
$\beta$ Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average 0 80.8 61.0 30.8 72.5 61.3 0.1 89.6 63.7 33.3 79.2 66.5 0.2 82.5 67.5 34.2 77.9 65.5 0.3 73.8 58.3 29.2 85.2 61.6
We analyze the impact of hyperparameters $\lambda$ and $\beta$ based on the results in Table 6 and Table 7. First, regarding the prior loss weight $\lambda$, we observe that even when $\lambda=0$ (where the prior branch is used solely for LLR calculation without explicit action supervision), the model achieves an average success rate of 63.3%. This represents a substantial improvement over the QwenGR00T baseline (55.2%), confirming that the LLR objective alone effectively regularizes the policy. Increasing $\lambda$ to 0.3 yields the optimal performance of 66.5%, validating that explicitly learning the vision-only prior $p(a|v)$ further aids the decomposition. The performance remains robust across $\lambda\in[0,0.5]$, though it drops slightly at higher values.
Second, for the LLR weight $\beta$, setting $\beta=0$ corresponds to training the dual-branch architecture without the mutual information maximization term. This configuration achieves 61.3%, surpassing the baseline (55.2%). This suggests that the architectural design, which explicitly separates a vision-only pathway to absorb dataset biases, inherently helps the posterior branch focus on language. Incorporating the LLR term ($\beta=0.1$) further boosts the success rate to 66.5%.
C.2 Ablation on Number of Latent Action Queries
To explore the potential of Latent Action Queries as a out-of-the-box technique, we conduct this ablation solely on the query mechanism, without incorporating our proposed Bayesian decomposition. We analyse the impact of the number of latent action queries on model performance, as summarized in Table 8. We observe that increasing the number of queries from 16 to 64 leads to a substantial improvement in success rate. However, doubling the queries to 128 results in performance saturation without providing additional benefits. Consequently, we adopt 64 queries as the optimal configuration to balance performance and computational overhead.
Table 8: Ablation study on Number of Latent Action Query. All experiments are based on the Qwen3-VL-4B backbone.
Number Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average 16 69.8 47.9 20.8 60.4 49.7 32 62.5 57.4 25.8 79.2 56.2 64 74.6 58.3 29.2 67.9 57.5 128 72.9 57.3 29.1 70.8 57.5